How many calories of heat need to be extracted from 200.g of steam at 120.°C to change it into ice at --12.0°C ?

1
Expert's answer
2022-03-06T15:17:19-0500

In order to cool down the steam to "100\degree C" the following amount of heat should be extracted:


"Q_1 =c_sm\Delta T"

where "c_s = 0.48cal/g/\degree C" is the specific heat capacity of steam, "m = 200g" is the mass of steam, and "\Delta T = 120\degree C - 100\degree C = 20\degree C". Thus, obtain:


"Q_1 = 0.48\cdot 200\cdot 20 = 1920cal"


In order to convert steem to water the following amount of heat should be extracted:


"Q_2 = Lm"

where "L = 540cal\/g" is the latent heat of vaporization of steam. Thus, obtain:

"Q_2 = 540\cdot 200 = 108000cal"


In order to cool down water from "100\degree C" to "0\degree C" the following amount of heat should be extracted:


"Q_3 = c_wm\Delta T"

where "c_w = 1.00cal/g/\degree C" is the specific heat capacity of water, and "\Delta T = 100\degree C". Thus, obtain:


"Q_3 = 1\cdot 200\cdot 100 = 20000cal"

In order co convert water to ice the following amount of heat should be extracted:


"Q_4 = \lambda m"

where "\lambda = 79.7cal/kg" is the enthalpy of fusion of water. Thus, obtain:


"Q_4 = 79.7\cdot 200 = 15940cal"

In order to cool down the ice to "-12\degree C" the following amount of heat should be extracted:


"Q_5 = c_im\Delta T"

where "c_i = 0.5cal/g/\degree C" is the specific heat capacity of ice, and "\Delta T = 12\degree C". Thus, obtain:


"Q_5 = 0.5\cdot 200 = 100cal"

The total heat is then:


"Q = Q_1 + Q_2+Q_3+Q_4+Q_5 = 145960cal\approx 1.46\times 10^5cal"

Answer. "1.46\times 10^5cal"